Geotail CPI Survey Plots

L. A. Frank, K. L. Ackerson, W. R. Paterson, R. L. Dyson, G. T. Parmentier, and G. L. Pickett

1 Summary

The Comprehensive Plasma Instrumentation (CPI) for the Geotail spacecraft consists of three plasma analyzers: (1) an electrostatic analyzer for hot electrons and ions, (2) a complementary electrostatic analyzer for cool plasmas with high bulk speeds such as those found within the solar wind and magnetosheath, and (3) an ion composition analyzer for identification of ion species such as H^{+}, He^{+}, He^{++}, and O^{+}. The energy-per-unit charge (E/Q) ranges for these analyzers are 1.3 V to 48.2 kV, 145 V to 6830 V, and 1.3 V to 48.2 kV, respectively. Three-dimensional velocity distributions of the ions and electrons are acquired for the first time in the distant magnetotail and are sufficiently accurate to separate E⃗ × B⃗ drifts from field-aligned flows associated with currents. Among the many specific scientific objectives of the plasma observations are the relative importance of ionospheric and solar wind sources, searches for plasmoids and flux ropes, current-sheet acceleration of plasmas, and convection and current systems in the magnetotail.

2 Description of Instrumentation

The Comprehensive Plasma Instrumentation is distributed into three separate mechanical housings. CPI-HS is the Hot Plasma Analyzer and the Solar Wind Ion Analyzer, and CPI-IC is the Ion Composition Analyzer. CPI-E contains the interface circuits to the spacecraft and the microprocessor responsible for overall control of the operation of the instrument. A simplified block diagram is included in Figure 1. The instrumentation is equipped with three Harris 80C85 microprocessors, each with read-only memory (ROM) and random-access memory (RAM) with capacities of 16 kilobytes and 48 kilobytes, respectively. The total mass and average power of the plasma instrumentation is 17.5 kg and 10.5 W, respectively. The telemetry allotment is 4.6 kilobits/s and is extended by runlength coding in the instrument to about 8.2 kilobits/s.

The configuration diagram for CPI-HS is shown in Figure 2. The Hot Plasma Analyzer is equipped with three spherical-segment plates. The outer and inner plates are grounded, and the center plate is supplied with a programmable series of voltages, 65 mV to 2.4 kV, in order to provide E/Q analyses of the ions and electrons. The radius of curvature R of the center plate is 10.2 cm and the analyzer constant, R/(2ΔR), is 20 for electrons and positive ions. The 64 E/Q passbands for the range 1.3V ≤ E/Q ≤ 48.2kV are identified in Table 1. The inner and outer channels are the positive ion and electron analyzers, respectively. Each of these two analyzers is serviced with 9 continuous-channel electron multipliers in order to divide the fan-shaped acceptance angle into 9 individual segments. These sensors are Galileo Electro-Optics Spiraltrons with either 1 mm or 3 mm diameter acceptance apertures, the outputs of which are fed into individual AMPTEK Inc. model A111F PAD amplifier-discriminators. The Spiraltron bias voltages are controlled by ground commands to any one of 32 voltage levels between +1690 to +3400 V (electron sensors) and −1718 to −3450 V (positive ion sensors). Considerable design effort was expended in shaping the analyzer fields-of-view and in maintaining approximately equal geometric factors and E/Q resolutions for the two sets of nine sensors. These goals were accomplished with a fence at the entrance apertures of the electrostatic analyzers, tailored angular length (analyzer arc angle) from entrance to exit apertures, and aperture plates at the sensors. In addition the concave surfaces of the spherical-segment plates were machined with sawtooth serrations of 1-mm depth in order to suppress the ultraviolet (primarily solar Lyman−α) and electron scattering along the plates into the sensors.

The fields-of-view, energy resolutions and provisional geometric factors are derived from ray tracing and laboratory calibrations and are tabulated in Table 2. The provisional geometric factors are to be refined with postlaunch data. The latitude θ is taken with respect to the spin axis of the spacecraft with 0^{∘} in a direction parallel to the spin axis and viewing generally toward the south ecliptic pole. The azimuthal range of ϕ is given in the spacecraft spherical coordinate system. The meridional plane at ϕ = 0^{∘} is parallel to the normal to the entrance aperture of the electrostatic analyzers. For example, the instantaneous field-of-view for electron sensor E1 is the solid angle within 4^{∘} ≤ θ ≤ 38^{∘} and 31^{∘} ≤ ϕ ≤ 81^{∘}. Note that the total coverage of latitude is 4^{∘} to 176^{∘} and allows sampling of all but 0.3% of the 4π−sr solid angle for particle velocity vectors at the spacecraft position. This is an important feature for accurate determination of the three-dimensional velocity distributions.

The operation of the Hot Plasma Analyzer and other two analyzers to be described below, is extremely flexible and is controlled by uplinked commands to the instrument microprocessors. A frequent operating mode for measuring the three-dimensional velocity distributions of positive ions and electrons is the division of one spacecraft rotation period (3 s) into eight equal intervals (azimuthal sectors). Four E/Q passbands are consecutively sampled in each of these azimuthal sectors. Every other passband in the range 22.2 V to 48.2 kV (see Table 1) is used. Thus a 1728-sample determination of each of the ion and electron intensities over the above E/Q range is acquired in 18 s. The repetition rate is 21 s because 2304 samples each of the ion and electron velocity distributions, 1.3 V ≤ E/Q ≤ 230 V, are gathered each 192 s in order to characterize the spacecraft potential, the secondary electron environment, and ambient cold plasmas.

The configuration of the Solar Wind Ion Analyzer is also shown in Figure 2. The twelve sets of sensors and amplifier-discriminators are the same types as those employed for the Hot Plasma Analyzer. The bias voltage for the Spiraltrons can be varied from −1975 V to −3940 V in 32 increments by ground command. The analyzer arc angle is 75^{∘} and the analyzer constant is 42. A programmed series of potentials ranging from 3.4 V to 163 V on the outer plate with the inner plate grounded provides measurements of positive ions with 144 V ≤ E/Q ≤ 6825 V within a maximum of 64 passbands. These passbands are listed in Table 3. The total latitudinal span of the field-of-view is 60^{∘} ≤ θ ≤ 120^{∘} which is divided into 12 equal, contiguous segments of 5^{∘} with the use of 12 sensors at the exit aperture of the electrostatic analyzer. The angular coverages, energy resolutions, and provisional geometric factors for the 12 sensors are tabulated in Table 4.

Operation of the Solar Wind Ion Analyzer is flexible via ground commands. A current operating mode provides division of the azimuthal (spin) sector ϕ_{s} = 294.5^{∘} to 90.0^{∘} into 42 equal, contiguous segments where ϕ_{s} = 0^{∘} is the direction to the sun. Thus the field-of-view consists of 504 contiguous solid angles, 12 (latitude) × 42 (azimuth). These fields-of-view and the E/Q passbands are interleaved to achieve 7560 samples of the ion velocity distribution once each 48 s. The odd-numbered passbands as listed in Table 3 are employed for this particular mode of operation.

A pictorial diagram of the Ion Composition Analyzer is shown in Figure 3. The radius of the inner plate is 11.3 cm and the plate separation is chosen to provide an analyzer constant of 20. This analyzer constant is the same as that for the Hot Plasma Analyser and allows easy implementation of beam search modes for the Ion Composition analyzer as determined by onboard processing of the ion responses of the Hot Plasma Analyzer. The E/Q range is 1.3 V to 48.2 kV and the 64 passbands are listed in Table 1. At the exit aperture of the electrostatic analyzer are positioned five imaging mass spectrometers which provide angular coverage of the mass-per-unit charge (M/Q) composition of positive ions.

A diagram of one of these mass spectrometers is shown in Figure 4. The first element is a slit collimator which confines the acceptance solid angle to 30^{∘} × 1^{∘} and is positioned at the exit aperture of the electrostatic analyzer. Dispersion of the positive ion beam is then provided by a small electromagnet. The pole pieces of the magnet are 0.4 cm^{2} in area with a 0.4-cm gap. The magnetic induction B in the gap is stepped from 0.005 T to 0.32 T via a programmed current supply. Extensive laboratory effort has verified the focusing capability of the combined gap and fringe magnetic fields, which is necessary in order to effectively apply this small magnet in a mass spectrometer. The dispersed ion beam is incident on a Galileo Electro-Optics MCP–25/12 microchannel plate. A ground screen is placed in front of the microchannel plate and the ions are post-accelerated into the microchannel plate by about −300 V to accommodate the detection of low-energy ions. The chevron-type microchannel plate has a 10-mm diameter, 25 μm pore size, and a saturated gain of ∼ 10^{6}.

In order to extend the effective operating lifetime of the microchannel plate, a supplementary gain is obtained by post-accelerating the output of the microchannel plate onto a metallized phosphor. The spacing between the exit aperture of the microchannel plate and phosphor is such that the output pulses of the microchannel plate are proximity-focused onto the phosphor. The phosphor is JEDEC P31 with a quantum yield factor of ∼ 0.05 photons/eV. Thus positive ions with a given M/Q pass through the slit collimator, are deflected by the magnet, produce electron pulses in the microchannel plate, and give a thin geometrical line of red light emission on the phosphor. The spatial resolution of the microchannel plate is thusly used to obtain a large number of mass channels, simultaneously and separately, over a large M/Q range.

The light from the phosphor is fed by tapered fiber optics to an EG&G Reticon photodiode array, Model RL0512SAF. Each of the 512 elements of this array is dimensioned 25 μm × 2.5 mm with the longer dimension aligned parallel to the above line of emission corresponding to a given M/Q species. The peak photodiode efficiency is ∼ 80% at 650 nm. The Reticon photodiode arrays are operated such that power- and mass-consuming refrigeration devices are unnecessary to achieve single-pulse (ion) sensitivity. The photodiode arrays are operated in flight at temperatures ∼ 0^{∘} by passive cooling with a radiator and at a 40 ms sampling period. The thermionic background is reduced to values below read-out threshold, and, if necessary, a low level of fixed-pattern electronic noise is subtracted by onboard processing. The individual photodiode saturation charge is ∼ 14pC. These photodiode analog outputs are linearly digitized into 256 bins (0 to 255 digitization units, dn). The gain of the microchannel plate and phosphor is controlled by ground command. The microchannel plate bias voltage and the phosphor bias voltage can be set to any one of 32 sets of values in the ranges 1790 V to 3060 V and 7300 V to 9100 V, respectively. Proximity focusing of the output of the microchannel plate onto the phosphor allows an additional gain margin of ∼ 50. That is, severe degradation of the microchannel plate can occur and the sensor system will remain operational. In addition the amplifier gain for the Reticon output charge pulse can be varied by ground command. The maximum event rate for detection of ions with given M/Q is typically ∼ 10^{4}/s.

For the Ion Composition Analyzer the analyzer arc angles, acceptance angles, energy resolutions, and provisional geometric factors are listed in Table 5. The Reticon site numbers for the centers of the M/Q passbands for H^{+}, He^{++}, He^{+}, and O^{+} are also given at the bottom of this table. Site 0 corresponds to no ion deflection, or gap magnetic field 0T. These values are listed for every fifth E/Q passband, as well as the gap field. The shape of the M/Q passband as a function of Reticon site number is shown in Figure 5. For deflections ≥ 50 sites, the shape of the passband is almost independent of the E/Q and M/Q of the ion.

Microprocessor control of the sampling of the Ion Composition Analyzer is relatively complex when compared to the other two analyzers and includes survey and beam search modes. The beam search modes utilize the responses of the Hot Plasma Analyzer to positive ions to identify the direction (θ, ϕ_{s}) and E/Q of peak responses where ϕ_{s} is the spin azimuthal angle of the spacecraft. The mass spectrometer with field-of-view nearest θ is then sampled at the corresponding values of ϕ_{s} and E/Q. During the survey mode ion spectrometer with field-of-view perpendicular to the spacecraft spin axis is sampled in four angular sectors of 26^{∘} which are centered generally sunward, dawnward, tailward and duskward. For each 26^{∘}-sector four 512-sample M/Q spectra are taken and summed in the instrument processor. These M/Q spectra are acquired at a fixed value of E/Q during one spacecraft spin period. A survey mode cycle typically consists of 24 E/Q passbands. the corresponding repetition time is 32 spin periods (96 s) and includes the beam search modes.

3 Description of the Solar Wind Ion Survey Spectrogram

The responses of the Solar Wind Ion Analyzer were averaged for about 50 minutes. For each of the 32 E/Q passbands, the responses of all twelve sensors in all sectors were scanned, and the peak response for each passband was determined. The logarithm (base 10) of each of these responses was color coded and displayed in the Energy-time spectrogram.

Velocity moments of these averaged distribution functions were integrated to obtain the ion number density and flow velocity:

n = ∫ fdv⃗ = number density

u⃗ = \frac{1}{n} ∫ fv⃗dv⃗ = flow velocity

The number density is plotted in the second panel and the magnitude of the flow velocity is shown in the third panel. The polar angles of the flow velocity, ϕ (ϕ = 180^{∘} is anti-sunward) and θ (θ = 90^{∘} is the S/C spin plane) are plotted in fourth panel.

The kinetic temperature was computed from the distribution function by integrating:

T = \frac{m}{3Kn} ∫ |u⃗ − v⃗|^{2}fdv⃗ = kinetic temperature

The CPI solar wind ion analyzer is an E/Q analyzer and the initial distribution functions were calculated assuming that the plasma is H^{+}. Care must be taken when multi-component plasmas are present. If He^{++} is present and is flowing with the same speed as the H^{+} component, it will contribute to the distribution function at twice the velocity (four times the energy) of the H^{+}. To facilitate the automated processing of these data the above integral was evaluated between |v⃗_{min}| = 171km/s and |v⃗_{max}| = 1.35 × |u⃗|. The resulting temperatures produce excellent fits to the observations. When He^{++} is present the distribution functions are consistent with an He^{++} temperature, T_{He^{++}} = 4 × T_{H^{+}}, i.e., the H^{+} and He^{++} have similar thermal velocities.

4 Description of the Hot Plasma Survey Plot

The Hot Plasma analyzer measures the intensities of electrically charged particles as functions of particle energy and direction. Measurements of electrons and positive ions are acquired simultaneously. This analyzer is one element of the Comprehensive Plasma Instrumentation (CPI) on board the Geotail spacecraft. Detailed descriptions of the Hot Plasma analyzer (HP), the Solar Wind analyzer (SW), the Ion Composition analyzer (IC), and the associated electronics package that comprise the Comprehensive Plasma Instrumentation are provided in the Description of Instrumentation.

Measurements from the Hot Plasma analyzer are available at this Web site in a graphical survey format. This format is designed to provide convenient access to the observations. Each survey plot is a five-day record of the electron and ion intensities and several parameters derived from those intensities. The measurements are given in a sequence of five panels. From top to bottom the panels display:

1. N [cm^{−3}], the ion number density.

2. V_{X} [km s^{−1}], the component of the ion bulk-flow velocity along the Earth-Sun line. Positive flow is towards the sun.

3. T_{i} and T_{e} [K], the respective temperatures of ions and electrons.

4. The intensities of positive ions as a function of particle energy.

5. The intensities of electrons as a function of particle energy.

The particle intensities given in panels 4 and 5 are five-minute accumulations from the ion and electron sensors that view the ecliptic plane. This plane is perpendicular to the spin axis of the spacecraft, and the intensities are summed over eight equally spaced spin sectors. The particle intensities are color coded according to the logarithmic color bar on the right-hand side of the figure which has separate scales for ions and electrons. Particle energies-per-charge (E/Q, Volts) are indicated by the logarithmic scale on the ordinates of the ion and electron panels. Dates corresponding to the major tick marks on the abscissa are shown at the bottom of the display. Minor tick marks are given at four-hour intervals. The projection of the Geotail orbit in the ecliptic plane is shown for reference at the top of the display. The beginning of each day is marked with a dot, and the first day of the five-day sequence is highlighted.

The plasma parameters given in the first three panels of the survey display are computed for particles with energies-per-charge in the range 50 V to 48 kV. The ions are assumed to be protons. The parameters are computed using five-minute accumulations from the electron and ion sensors that view the ecliptic plane. Thus, the number density, the V_{X} component of the flow, and the temperatures given here are equivalent to those that would be gained with a two-dimensional plasma analyzer.

5 Acknowledgments

The authors extremely appreciate the efforts of D. W. Burke, P. K. Hardebeck, J. D. Harvey, L. M. Henry, K. A. Hunt, P. S. Reilly, and J. A. Sartain from The University of Iowa, and M. H. Acuña, J. P. Galloway, D. E. Galoppo, M. M. Grant, and K. O. Sizemore and many other persons of the Goddard Space Flight Center/NASA in helping us implement this investigation. Invaluable assistance and great courtesy were given to us in prelaunch integration and testing and in postlaunch operations of our instrument by H. Hayakawa, S. Machida, T. Mukai, M. Nakamura, A. Nishida, T. Obara, K. Tsuruda, and K. Uesugi at the Institute of Space and Astronautical Science, J. Ishida of NEC, and a large number of their colleagues. For a more complete discussion of the Comprehensive Plasma Instrumentation for the Geotail spacecraft see J. Geomag. Geoelectr., 46, 23–37, 1994.

6 Tables

Table 1. Hot Plasma Analyzer and Ion Composition Analyzer Center of E/Q Passbands

Passband            Passband             Passband
No.     E/Q         No.     E/Q          No.     E/Q
--------  ------    --------  ------     --------   -------
 0      1.3 V       22     51.2 V        43      1.71 kV
 1      1.5         23     60.5          44      2.02
 2      1.8         24     71.5          45      2.39
 3      2.1         25     84.5          46      2.82
 4      2.5         26     100           47      3.33
 5      3.0         27     118           48      3.94
 6      3.5         28     140           49      4.65
 7      4.2         29     165           50      5.50
 8      4.9         30     195           51      6.50
 9      5.8         31     230           52      7.68
10      6.9         32     322           53      9.08
11      8.2         33     322           54     10.7
12      9.6         34     380           55     12.7
13     11.4         35     449           56     15.0
14     13.5         36     531           57     17.7
15     15.9         37     627           58     20.9
16     18.8         38     741           59     24.7
17     22.2         39     876           60     29.2
18     26.2         40     1.04 kV       61     34.5
19     31.0         41     1.22          62     40.8
20     36.7         42     1.45          63     48.2
21     43.3

Table 2. Hot Plasma Analyzer Nominal Performance Parameters

      Analyzer        Angular           Energy         Geometric
      arc angle       Coverage         Resolution       Factor
                     --------------
Sensor  (deg.)        theta   phi (deg.)   dE/E       cm2-sr-eV/eV
--------------      -----------------------------------------------
E1       75          4, 38  -31, 81       0.10       8.0 x 10^{-6}
E2       70         22, 48  -20, 45       0.13       1.8 x 10^{-5}
E3       70         42, 66  -33, 23       0.08       2.1 x 10^{-5}
E4       65         57, 84  -22, 11       0.09       2.0 x 10^{-5}
E5       60         74,106  -12,  5       0.10       1.8 x 10^{-5}
E6       65         96,123  -22, 11       0.09       2.0 x 10^{-5}
E7       70        114,138  -33, 23       0.08       2.1 x 10^{-5}
E8       70        132,158  -20, 45       0.13       1.8 x 10^{-5}
E9       75        142,176  -31, 81       0.10       8.0 x 10^{-6}
P1       75          4, 35  -66, 81       0.09       4.8 x 10^{-5}
P2       70         22, 48  -19, 44       0.12       5.9 x 10^{-5}
P3       70         42, 66  -33, 22       0.08       5.9 x 10^{-5}
P4       65         57, 84  -22, 10       0.09       6.5 x 10^{-5}
P5       60         74,106  -12,  4       0.10       6.5 x 10^{-5}
P6       65         96,123  -22, 10       0.09       6.5 x 10^{-5}
P7       70        114,138  -33, 22       0.08       5.9 x 10^{-5}
P8       70        132,158  -19, 44       0.12       5.9 x 10^{-5}
P9       75        145,176  -66, 81       0.09       4.8 x 10^{-5}

Table 3. Solar Wind Ion Analyzer Center of E/Q Passbands

Passband            Passband             Passband
No.     E/Q         No.    E/Q           No.     E/Q
-------   -----     -------- ------      --------  ------
 0     144 V        22     555 V         43     2010 V
 1     153          23     590           44     2133
 2     163          24     627           45     2270
 3     173          25     667           46     2410
 4     184          26     709           47     2560
 5     196          27     753           48     2725
 6     208          28     801           49     2900
 7     221          29     852           50     3080
 8     235          30     905           51     3270
 9     250          31     962           52     3480
10     266          32    1020           53     3700
11     283          33    1090           54     3930
12     301          34    1160           55     4180
13     320          35    1230           56     4450
14     340          36    1310           57     4730
15     361          37    1390           58     5025
16     384          38    1480           59     5340
17     408          39    1570           60     5680
18     434          40    1670           61     6040
19     462          41    1775           62     6420
20     491          42    1890           63     6825
21     522

Table 4. Solar Wind Ion Analyzer Nominal Performance Parameters

                 Angular           Energy        Geometric
                Coverage         Resolution        Factor
               -----------
Sensor          theta   phi (deg.)   dE/E        cm2-sr-eV/eV
------         ------------------------------------------------
S1             60,66  -8,7          0.034       3.3 x 10^{-6}
S2             65,71  -7,6          0.034       3.5 x 10^{-6}
S3             70,76  -6,5          0.034       3.6 x 10^{-6}
S4             75,81  -5,4          0.034       3.6 x 10^{-6}
S5             79,86  -4,3          0.034       3.7 x 10^{-6}
S6             84,91  -3,2          0.034       3.7 x 10^{-6}
S7             89,96  -3,2          0.034       3.7 x 10^{-6}
S8             94,101 -4,3          0.034       3.7 x 10^{-6}
S9             99,105 -5,4          0.034       3.6 x 10^{-6}
S10           104,110 -6,5          0.034       3.6 x 10^{-6}
S11           109,115 -7,6          0.034       3.5 x 10^{-6}
S12           114,120 -8,7          0.034       3.3 x 10^{-6}